OpenStudy (anonymous):

Currently, Sam's coach is twice as old as Sam. When Sam started soccer 10 years ago, his coach was 3 times his age. How old is Sam now?

OpenStudy (anonymous):

C = 2S (C-10) = 3(S-10) C = 40 S = 20